# ASPD: Unlocking Adaptive Serial-Parallel Decoding by Exploring Intrinsic Parallelism in LLMs

### v1

@@ 4.5 Ablation Study
While autoregressive models inherently possess parallelizable characteristics in their responses, precisely identifying parallelizable components remains a significant challenge. As discussed in Section 3.1, we must ensure semantic-level parallelization while maintaining branch independence and completeness. Table 8 demonstrates the performance variations across different data processing methodologies. For our comparative analysis, we implemented APAR(Liu et al., 2024) using their official codebase and PASTA(Jin et al., 2025) using the prompt provided in their paper (as their official code is not open-sourced). All construction processes utilized the Qwen3-235B-22A(Yang et al., 2025) model.
@@
During parallel decoding, multiple branches decode simultaneously while remaining mutually invisible, making the arrangement of position encodings particularly crucial. We abstract position encoding schemes into the following categories:
Parallel Branches with Pre-allocated Fixed-Length Position IDs. Pre-allocated fixed-length position encoding intervals for each parallel branch, which is adopted by PASTA (Jin et al., 2025). Our experiments show that with fixed intervals, model performance peaks at length 32 (score 6.09) but degrades significantly with larger intervals (4.24 at 512), suggesting an optimal interval matching the average parallel branch length in training data.
Parallel Branches with Same Position IDs. All parallel branches share identical position encodings at each timestamp, with three variants:
  • Same-Max: Uses maximum position encoding across parallel branches for merging, achieving good performance (6.78) with moderate speed (89.45).
  • Same-Rearrange(Same-Re): Reorders position encodings sequentially post-decoding, requiring KV-cache re-prefill which introduces additional overhead.This implementation achieves suboptimal performance (7.29) and speed (95.24), representing a compromise between computational efficiency and generation quality.
  • Same-Sequential(Same-Seq): Assigns actual sequential positions, demonstrating optimal balance between performance (7.64) and speed (104.21).The performance degradation of Same-Re relative to Same-Seq can be attributed to the inconsistency between position ids in the parallel phase and subsequent sequential phase during training. This discrepancy likely disrupts the model’s optimization trajectory, ultimately resulting in suboptimal performance metrics.
@@ 5 Conclusion
In this work, we present ASPD: The Adaptive Serial-Paralle Decoding framework for efficient hybird decoding within auto regressive large language models. The proposed method introduces a non-invasive parallel data transformation pipeline and internal parallelization with branch-invisible attention masks and shared position encodings, achieving substantial latency reduction while maintaining response quality compared to traditional autoregressive LLMs.

### v2

@@ 4.5 Ablation Study
While autoregressive models inherently possess parallelizable characteristics in their responses, precisely identifying parallelizable components remains a significant challenge. As discussed in Section 3.1, we must ensure semantic-level parallelization while maintaining branch independence and completeness. Table 8 demonstrates the performance variations across different data processing methodologies. For our comparative analysis, we implemented APAR(Liu et al., 2024) using their official codebase and PASTA(Jin et al., 2025) using the prompt provided in their paper (as their official code is not open-sourced). All construction processes utilized the Qwen3-235B-22A(Yang et al., 2025b) model.
@@
During parallel decoding, multiple branches decode simultaneously while remaining mutually invisible, making the arrangement of position IDs particularly crucial. We abstract position encoding schemes into the following categories:
Parallel Branches with Pre-allocated Fixed-Length Position IDs. Pre-allocated fixed-length position ID intervals for each parallel branch, which are adopted by PASTA (Jin et al., 2025). Our experiments show that with fixed intervals, model performance peaks at length 32 (score 6.09) but degrades significantly with larger intervals (4.24 at 512), suggesting an optimal interval matching the average parallel branch length in training data.
Parallel Branches with Same Position IDs. All parallel branches share identical position ID at each timestamp, with three variants:
  • Same-Max: Uses maximum position ID across parallel branches for merging, achieving good performance (6.78) with moderate speed (89.45).
  • Same-Rearrange(Same-Re): Reorders position IDs sequentially post-decoding, requiring KV-cache re-prefill which introduces additional overhead.This implementation achieves suboptimal performance (7.29) and speed (95.24), representing a compromise between computational efficiency and generation quality.
  • Same-Sequential(Same-Seq): Assigns actual sequential position IDs, demonstrating optimal balance between performance (7.64) and speed (104.21).The performance degradation of Same-Re relative to Same-Seq can be attributed to the inconsistency between position IDs in the parallel phase and subsequent sequential phase during training. This discrepancy likely disrupts the model’s optimization trajectory, ultimately resulting in suboptimal performance metrics.
@@ 5 Conclusion
In this work, we present ASPD: The Adaptive Serial-Paralle Decoding framework for efficient hybird decoding within auto regressive large language models. The proposed method introduces a non-invasive parallel data transformation pipeline and internal parallelization with branch-invisible attention masks and shared position IDs, achieving substantial latency reduction while maintaining response quality compared to traditional autoregressive LLMs.